OpenStudy (anonymous):

Find the inclination θ (in radians AND degrees) of a line passing through the points: (-7, -1) (-1, 2)

OpenStudy (anonymous):

find the slope which is tan theta..

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

(2--1)/(-1--7)=slope=tan theta

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (anonymous):

no..You can get it from this

OpenStudy (anonymous):

0.464 radian(s) or 26.565° -0.464 radian(s) or -26.565° -2.678 radian(s) or 153.435° 1.107 radian(s) or 63.435° -1.107 radian(s) or -63.435°

OpenStudy (anonymous):

answer choices^^

OpenStudy (anonymous):

ok..ok

OpenStudy (anonymous):

actually the slope was 3/6=1/2

OpenStudy (anonymous):

answer is 26.565 degrees

OpenStudy (anonymous):

@lamp12345 it was just a miscalculation...it is 26.565 degrees

OpenStudy (anonymous):

I will just delete the wrong ones..:)

OpenStudy (anonymous):

got it??